OpenStudy (anonymous):

The lengths of the sides of a rectangle become 3 times greater than the original measurements. Which statement is true about the rectangle? Its area becomes 3 times greater, and perimeter becomes 9 times greater. Its area becomes 6 times greater, and perimeter becomes 3 times greater. Its area becomes 3 times greater, and perimeter becomes 6 times greater. Its area becomes 9 times greater, and perimeter becomes 3 times greater.

Parth (parthkohli):

There's a trick for this. Think of length as 2 and width as 1. The area will be 2 * 1 = 2 and perimeter would be 2 + 2 + 1 + 1 = 6 Now make the length as 6 and width as 3. The area would be 6 * 3 = 18 and perimeter would be 6 + 6 + 3 + 3 = 18.

Parth (parthkohli):

Perimeters: 6 becomes 18 Areas : 2 becomes 18

OpenStudy (anonymous):

so which one wou.d be the answer would it be the seond one

Parth (parthkohli):

I see that the area is becoming 9 times larger.

OpenStudy (anonymous):

so the last one ??

Parth (parthkohli):

yep

OpenStudy (anonymous):

thank u